# Take It, Leave It, or Fix It: Measuring Productivity and Trust in Human-AI Collaboration

### v1

@@ Abstract.
Although recent developments in generative AI have greatly enhanced the capabilities of conversational agents such as Google’s Bard or OpenAI’s ChatGPT, it’s unclear whether the usage of these agents aids users across various contexts. To better understand how access to conversational AI affects productivity and trust, we conducted a mixed-methods, task-based user study, observing 76 software engineers (N=76) as they completed a programming exam with and without access to Bard. Effects on performance, efficiency, satisfaction, and trust vary depending on user expertise, question type (open-ended ”solve” questions vs. definitive ”search” questions), and measurement type (demonstrated vs. self-reported). Our findings include evidence of automation complacency, increased reliance on the AI over the course of the task, and increased performance for novices on “solve”-type questions when using the AI. We discuss common behaviors, design recommendations, and impact considerations to improve collaborations with conversational AI.
@@ 1. Introduction
Recent advancements in generative artificial intelligence (AI) have the potential to enhance productivity across domains such as medicine (Li et al., 2024), research (Salvagno et al., 2023; Lund and Wang, 2023), and technology (Rice et al., 2024). In the context of software development, conversational AI such as Google’s Bard and Open AI’s ChatGPT can generate code, and Github’s Copilot can autocomplete code. However, it’s unclear whether these systems strictly improve productivity. State-of-the-art agents suffer from imperfect accuracy and biases (Oviedo-Trespalacios et al., 2023; Wach et al., 2023), and humans have demonstrated cognitive biases such as automation bias and effort substitution when using LLM-based systems such as ChatGPT and Copilot (Al Madi, 2023; Barke et al., 2023; Vaithilingam et al., 2022; Noy and Zhang, 2023). Furthermore, behaviors and outcomes may depend on the expertise or literacy of the user; novices have been found to be less discerning of automation errors (Oviedo-Trespalacios et al., 2023).
@@ 3.2. Task design
Figure 1. First pass on a Bard-first, solve-type question.
@@
Figure 2. Second pass on a Bard-first, solve-type question.
@@ 4.1. Productivity
Figure 3. A structural equation model showing correlations between expertise measures and prodcutivity outcomes; $\beta$ is the normalized effect size in standard deviations, and $S$ denotes standard error.${}^{4}$
@@
Figure 3. A structural equation model showing correlations between expertise measures and prodcutivity outcomes; $\beta$ is the normalized effect size in standard deviations, and $S$ denotes standard error.${}^{4}$
@@ 4.1.2. Efficiency
Table 1. Left, a) Descriptive statistics on efficiency. Mean time spent per section in minutes, with sd. in parentheses. The Ind. column displays significance values from a two-sample independence t-test between Bard and Book times. Right, b) Regression coefficients and standard errors from a regression of time spent (in minutes) on expertise percentile rank.${}^{5}$
@@ 5. I feel more satisfied with my work completing
Table 2. Comparative satisfaction survey (Ziegler et al., 2022) administered post-task.${}^{6}$
@@ 7 Book Bard
Figure 5. Answer-changing percentages between passes.
@@ 2. Effect of expressing distrust on performance
Solve-type questions
0.09
@@
Search-type questions
-0.12
@@ Acknowledgements.
We thank our pilot and study participants, Google’s People and AI Research (PAIR) team, and the MIT Economics department for their valuable advice and support, notably Michael Terry, Lucas Dixon, Carrie Cai, Tobias Salz, Nikhil Agarwal, Ryan Mullins, Robert Upton, Martin Wattenberg, Fernanda Viégas, Hal Ableson, Mike Schaekermann, Key Lee, Andrea Knight Dolan, Mahima Pushkarna, and Hansa Srinivasan.

### v2

@@ Abstract.
Although recent developments in generative AI have greatly enhanced the capabilities of conversational agents such as Google’s Gemini or OpenAI’s ChatGPT, it’s unclear whether the usage of these agents aids users across various contexts. To better understand how access to conversational AI affects productivity and trust, we conducted a mixed-methods, task-based user study, observing 76 software engineers (N=76) as they completed a programming exam with and without access to Google’s Bard. [^1] At the time of this study in July 2023, Google’s generative artificial intelligence chatbot was named Bard. This has since been renamed to Gemini (Hsiao, 2024). Effects on performance, efficiency, satisfaction, and trust vary depending on user expertise, question type (open-ended ”solve” vs. definitive ”search” questions), and measurement type (demonstrated vs. self-reported). Our findings include evidence of automation complacency, increased reliance on the AI over the course of the task, and increased performance for novices on “solve”-type questions when using the AI. We discuss common behaviors, design recommendations, and impact considerations to improve collaborations with conversational AI.
large language models; user behaviors; artificial intelligence; trust
@@ 1. Introduction
Recent advancements in generative artificial intelligence (AI) have the potential to enhance productivity across domains such as medicine (Li et al., 2024), research (Salvagno et al., 2023; Lund and Wang, 2023), and technology (Rice et al., 2024). In the context of software development, conversational AI such as Google’s Gemini and Open AI’s ChatGPT can generate code, and Github’s Copilot can autocomplete code. However, it’s unclear whether these systems strictly improve productivity. State-of-the-art agents suffer from imperfect accuracy and biases (Oviedo-Trespalacios et al., 2023; Wach et al., 2023), and humans have demonstrated cognitive biases such as automation bias and effort substitution when using LLM-based systems such as ChatGPT and Copilot (Al Madi, 2023; Barke et al., 2023; Vaithilingam et al., 2022; Noy and Zhang, 2023). Furthermore, behaviors and outcomes may depend on the expertise or literacy of the user; novices have been found to be less discerning of automation errors (Oviedo-Trespalacios et al., 2023).
@@ 4.1. Productivity
Figure 3. A structural equation model showing correlations between expertise measures and prodcutivity outcomes; $\beta$ is the normalized effect size in standard deviations, and $S$ denotes standard error.4
[…]
@@ 4.1.2. Efficiency
Table 1. Left, a) Descriptive statistics on efficiency. Mean time spent per section in minutes, with sd. in parentheses. The Ind. column displays significance values from a two-sample independence t-test between Bard and Book times. Right, b) Regression coefficients and standard errors from a regression of time spent (in minutes) on expertise percentile rank.5
@@ 5. I feel more satisfied with my work completing
Table 2. Comparative satisfaction survey (Ziegler et al., 2022) administered post-task.6
@@ 2. Effect of expressing distrust on performance
Solve-type questions 0.09
@@
Search-type questions -0.12
@@ Acknowledgements.
We thank our pilot and study participants, Google’s People and AI Research (PAIR) team, and the MIT Economics department for their valuable advice and support, notably Michael Terry, Lucas Dixon, Carrie Cai, Tobias Salz, Nikhil Agarwal, Ryan Mullins, Robert Upton, Martin Wattenberg, Fernanda Viégas, Hal Ableson, Mike Schaekermann, Key Lee, Frank Schilbach, Andrea Knight Dolan, Mahima Pushkarna, and Hansa Srinivasan.